The stack on the left is made up of 15 pennies, and the stack on the right is also made up of 15 pennies. If the volume of the stack of pennies on the left is 390 mm3, what is the volume of the stack of pennies on the right in cubic millimeters? (Hint: only enter numerals in the answer blank) (4 points)  


1
Expert's answer
2021-10-19T01:57:43-0400

The volume of the stack of pennies on the right is also 390 cubic millimeters.

Proof

Given:- Volume of the stack of pennies on the left"= 390 mm^3\space 

\& \pi = 3.14"

now we know that volume "= \frac{4}{3}\space \pi (r^3)"

so

"\frac{4}{3}\space \pi (r^3) = 390 mm^3\\

r^3 =\frac{(390\times 3)}{4 }\times \pi\\

r^3 = 93.153"

"r = 4.53" (radius of a penny)

now for right stack volume :-

volume "= \frac{4}{3}\space \pi (r^3)"

where "r = 4.53 \space \& \pi = 3.14"

Volume of the stack of pennies on the right

"=\frac{ 4}{3} (3.14) ((4.53)^3)\\

= \frac{4}{3} (3.14) (93.153)\\

= 390 mm^3"

(proved)